OpenStudy (anonymous):

Verify the equation below with each of the values listed for t to find a solution. 4 = 3t – 2

OpenStudy (anonymous):

3*2=6-2=4 so t=2

OpenStudy (anonymous):

thnx @raynebow<3

OpenStudy (anonymous):

No problamo:)